OpenStudy (anonymous):

lim approaches x==> -2

OpenStudy (anonymous):

\[\sqrt[3]{\frac{ x+2 }{ x^{3}+8 }}\]

OpenStudy (welshfella):

if you plug in x = -2 you get the result 0/0 so I would try using L'hopital's rule

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\sqrt[3]\frac{x+2}{x^3+8}=\sqrt[3]\frac{x+2}{(x+2)(x^2-2x+4)}=\frac{1}{\sqrt[3]{x^2-2x+4}}\]

OpenStudy (welshfella):

thats the way

OpenStudy (anonymous):

thanks :D

OpenStudy (welshfella):

just plug in x = -2 and you have it